You are told that P(A)=.20, P(B)=.50 and the two events are independent. What is the probability of P(A or B)
Answer by jim_thompson5910(35256)

" the two events are independent", so

P(A and B) = P(A)*P(B)

P(A and B) = 0.20*0.50

P(A and B) = 0.10

-------------------------------------------------------

Now use this to find the probability of P(A or B)


P(A or B) = P(A) + P(B) - P(A and B)

P(A or B) = 0.20 + 0.50 - 0.10

P(A or B) = 0.70 - 0.10

P(A or B) = 0.60